OpenStudy (anonymous):

I really need help solving this. 0.5x-7<1.5+0.8x. Write in interval notation.

OpenStudy (jamesj):

First get all of the x terms to one side of the inequality. What do you get?

OpenStudy (anonymous):

-7<1.5+0.3x

OpenStudy (jamesj):

Yes. Now get x by itself on one side

OpenStudy (anonymous):

Ok so would it be -5.5/0.3<x?

OpenStudy (jamesj):

No. If you subtract 1.5 from both sides, you have 0.3x > -8.5 hence x > 85/3 = 28 1/3

OpenStudy (anonymous):

Oh ok. So how would I write that in interval notation?

OpenStudy (jamesj):

Well, if x > a for some constant a, how do you write that in interval notation? It is \[ x \in (a, \infty) \]

OpenStudy (anonymous):

so the answer would be (28,oo)? Im still confused.

OpenStudy (jamesj):

No. It is \[ x \in (28 \frac{1}{3}, \infty) \] or \[ x \in (83/3, \infty) \] whichever you prefer.

OpenStudy (anonymous):

oh ok. thanks you so much.